# Causal Graph Neural Networks for Healthcare

### v2

@@ 6.3 From methods to clinical practice
Box 2: Example of multi-modal validation framework The amyloid-$\beta$ hypothesis in Alzheimer’s disease Consider a causal GNN identifying a directed edge from amyloid-$\beta$ deposition to cognitive decline in Alzheimer’s disease, a relationship with validation complexities. Evidence triangulation: 1. Biological plausibility ($S_{\text{bio}}=+1$): Strong mechanistic support through the amyloid cascade hypothesis and decades of foundational neuroscience research. 2. Replication ($S_{\text{rep}}=+1$): Association replicates consistently across major cohorts, including ADNI, AIBL, and European EADC datasets with stable effect sizes. 3. Natural experiments ($S_{\text{quasi}}=-1$): Mendelian randomization studies using genetic variants affecting amyloid production demonstrate no causal effect on cognitive decline [93]. 4. Prospective validation ($S_{\text{prosp}}=-1$): Multiple Phase III clinical trials of amyloid-reducing therapies show minimal cognitive benefit despite substantial plaque reduction [94]. 5. Sensitivity analysis ($S_{\text{sens}}=0$): Moderate robustness to unmeasured confounding (E-value = 1.8) but substantial uncertainty regarding alternative causal structures. Aggregate evidence score: $\text{AES}=0.20(+1)+0.25(+1)+0.30(-1)+0.15(-1)+0.10(0)=0.00$ (4) Interpretation: Despite strong observational associations and biological plausibility, the null aggregate score reflects fundamentally contradictory evidence from high-quality natural experiments and prospective trials [83, 95]. This suggests the observational causal model may be confounded, with amyloid deposition potentially representing a biomarker of disease progression rather than a causal driver, or causality involving more complex dynamics, including threshold effects or critical developmental periods that the learned graph structure fails to capture. Clinical implication: Predictions based on this causal edge should be treated with appropriate caution, and amyloid reduction alone may not constitute an effective therapeutic strategy. This exemplifies how systematic evidence triangulation prevents overconfident causal claims despite compelling observational evidence, ultimately protecting patients from ineffective interventions whilst guiding research toward more productive therapeutic targets.
@@ 7 Conclusion
The transition from correlation-based prediction to mechanism-based medicine will define which artificial intelligence systems earn clinical trust and achieve sustainable deployment. This transformation requires balancing theoretical ambition with empirical humility, acknowledging that we construct pragmatic causal approximations that support clinical inferences at chosen levels of abstraction rather than discovering absolute causal truth in biological systems of extraordinary complexity. It demands computational sophistication matched with clinical interpretability, ensuring that mechanistic insights align with established biomedical knowledge and can be verified by domain experts. Most critically, it necessitates a transformative vision tempered by uncompromising validation standards, where novel causal discoveries undergo rigorous multi-modal triangulation before informing high-stakes clinical decisions. The methodological foundations are established through theoretical advances connecting structural causal models to differentiable graph architectures, empirical demonstrations across diverse healthcare applications, and emerging frameworks for validation under observational constraints [41, 15, 16]. The path forward demands not only technical innovation, but sustained collaboration between computational scientists developing causal inference methods, clinicians defining requirements and validation criteria, biomedical researchers providing mechanistic knowledge, and policymakers establishing appropriate oversight. The convergence of multi-modal biomedical data accumulation, regulatory demands for explainable artificial intelligence, and documented failures of correlation-based models under distribution shift creates both opportunity and urgency. Translation to clinical practice would enable healthcare systems to identify causal mechanisms rather than mere associations, predict interventional outcomes rather than observational patterns, and generate patient-specific counterfactuals that enable truly personalised medicine. This vision warrants continued investment provided that we maintain rigorous standards for genuine causal inference in complex biological systems, distinguish mechanistic understanding from pattern recognition, and prioritise patient outcomes over algorithmic sophistication.

### v3

@@ 6.4 From methods to clinical practice
Box 3: Example of multi-modal validation framework The amyloid-$\beta$ hypothesis in Alzheimer’s disease Consider a causal GNN identifying a directed edge from amyloid-$\beta$ deposition to cognitive decline in Alzheimer’s disease, a relationship with validation complexities. Evidence triangulation: 1. Biological plausibility ($S_{\text{bio}}=+1$): Strong mechanistic support through the amyloid cascade hypothesis and decades of foundational neuroscience research. 2. Replication ($S_{\text{rep}}=+1$): Association replicates consistently across major cohorts, including ADNI, AIBL, and European EADC datasets with stable effect sizes. 3. Natural experiments ($S_{\text{quasi}}=-1$): Mendelian randomization studies using genetic variants affecting amyloid production demonstrate no causal effect on cognitive decline [103]. 4. Prospective validation ($S_{\text{prosp}}=-1$): Multiple Phase III clinical trials of amyloid-reducing therapies show minimal cognitive benefit despite substantial plaque reduction [104]. 5. Sensitivity analysis ($S_{\text{sens}}=0$): Moderate robustness to unmeasured confounding (E-value = 1.8) but substantial uncertainty regarding alternative causal structures. Aggregate evidence score: $\text{AES}=0.20(+1)+0.25(+1)+0.30(-1)+0.15(-1)+0.10(0)=0.00$ (4) Interpretation: Despite strong observational associations and biological plausibility, the null aggregate score reflects fundamentally contradictory evidence from high-quality natural experiments and prospective trials [97, 105]. This suggests the observational causal model may be confounded, with amyloid deposition potentially representing a biomarker of disease progression rather than a causal driver, or causality involving more complex dynamics, including threshold effects or critical developmental periods that the learned graph structure fails to capture. Clinical implication: Predictions based on this causal edge should be treated with appropriate caution, and amyloid reduction alone may not constitute an effective therapeutic strategy. This exemplifies how systematic evidence triangulation prevents overconfident causal claims despite compelling observational evidence, ultimately protecting patients from ineffective interventions whilst guiding research toward more productive therapeutic targets.
[…]
We note that healthcare presents additional methodological challenges where causal reasoning proves essential but which exceed the scope of this Perspective. Small-sample learning remains problematic for causal GNNs, as deep learning models require large datasets to avoid overfitting while neuroimaging studies typically involve high-dimensional graphs with limited patient cohorts [107]. Multi-centre generalisation, though addressed by invariant learning (Section 3.4), involves domain adaptation complexities requiring specialised harmonisation techniques beyond causal invariance alone [108]. Time-varying confounding in longitudinal treatment regimes demands methods such as marginal structural models and g-computation that extend beyond the cross-sectional causal graphs emphasised here [109]. Missing data mechanisms, ubiquitous in electronic health records, require causal models for principled imputation yet remain inadequately addressed in current causal GNN frameworks [110]. These directions warrant dedicated methodological attention as the field matures.
@@ 7 Conclusion
The transition from correlation-based prediction to mechanism-based medicine will define which artificial intelligence systems earn clinical trust and achieve sustainable deployment. This transformation requires balancing theoretical ambition with empirical humility, acknowledging that we construct pragmatic causal approximations that support clinical inferences at chosen levels of abstraction rather than discovering absolute causal truth in biological systems of extraordinary complexity. It demands computational sophistication matched with clinical interpretability, ensuring that mechanistic insights align with established biomedical knowledge and can be verified by domain experts. Most critically, it necessitates a transformative vision tempered by uncompromising validation standards, where novel causal discoveries undergo rigorous multi-modal triangulation before informing high-stakes clinical decisions. The methodological foundations are established through theoretical advances connecting structural causal models to differentiable graph architectures, empirical demonstrations across diverse healthcare applications, and emerging frameworks for validation under observational constraints [45, 15, 16]. The path forward demands not only technical innovation, but sustained collaboration between computational scientists developing causal inference methods, clinicians defining requirements and validation criteria, biomedical researchers providing mechanistic knowledge, and policymakers establishing appropriate oversight. Multi-modal biomedical data accumulation, regulatory demands for explainable artificial intelligence, and documented failures of correlation-based models under distribution shift together create both opportunity and urgency. Translation to clinical practice would enable healthcare systems to identify causal mechanisms rather than mere associations, predict interventional outcomes rather than observational patterns, and generate patient-specific counterfactuals that enable truly personalised medicine. This vision warrants continued investment provided that we maintain rigorous standards for genuine causal inference in complex biological systems, distinguish mechanistic understanding from pattern recognition, and prioritise patient outcomes over algorithmic sophistication.